# Beyond Standard MoE: Mixture of Latent Experts for Resource-Efficient Language Models

### v1

@@ 4 Theoretical Analysis of MoE to MoLE Transformation
In this section, we establish the theoretical foundation for transforming a standard Mixture of Experts (MoE) model into its corresponding Mixture of Latent Experts (MoLE) counterpart. We focus on the “up operator” as a representative example. For analytical clarity, we make two simplifying assumptions: (1) we omit the subscript of $W_{\text{up}}^{i}$ for ease of notation during this analysis, and (2) we consider the case where $k=N$, implying a single shared latent space operator for all experts.
Keep the “down operator”. Note that for the “down operator” $W_{\text{down}}$, it maps the intermediate term from a low dimension into a high dimension, which captures the intrinsic properties of the given expert. Moreover, empirical experiments in Appendix C show that maintaining the structure of the “down operator” performs much better than not. Hence, we maintain the “down operator” in the transformed MoLE architecture and only change the “up operator” and “gate operator” from the MoE structure into the MoLE structure.
@@ 4.1 Transformation via Matrix Factorization
For a given weight matrix $W^{i}$ associated with expert $i$, we aim to find corresponding matrices $A^{i}$ and $B$ such that $W^{i}x\approx A^{i}Bx$ for any input $x$. This transformation represents the post-training perspective where we seek to convert pre-trained MoE parameters into the MoLE architecture.
@@
$\min_{A^{i},B}\quad F(A^{1},\cdots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}-A^{i}B\|_{F}^{2},$ (4.1)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. Problem equation 4.1 admits the theoretical optimal solutions by using the SVD decomposition technique.
Closed-Form Solution via Singular Value Decomposition. Obviously, problem equation 4.1 has infinitely many solutions since one can change one optimal solution $(A^{*},B^{*})$ into another by selecting a constant $\lambda$ and obtain another optimal solution $(\lambda A^{*},\frac{1}{\lambda}B^{*})$. Hence, we only provide “one” optimal solution here.
To derive a closed-form solution to problem equation 4.1, we consolidate the matrices into the following block structures:
$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix}.$ (4.2)
With this notation, problem equation 4.1 can be reformulated as:
$\min_{A,B}\quad\frac{1}{2}\|W-AB\|_{F}^{2},$ (4.3)
@@
While problem equation 4.3 has infinitely many solutions due to its underdetermined nature, the Eckart-Young-Mirsky theorem (Schmidt, 1989) provides an optimal solution with respect to the Frobenius norm. Specifically, let $W=U\Sigma V^{\top}$ be the singular value decomposition (SVD) of $W$, where:
@@
Let $\Sigma_{m}$ be the truncated version of $\Sigma$ that retains only the $m$ largest singular values: $\Sigma_{m}=\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m},0,\ldots,0)\in\mathbb{R}^{mN\times n}$. According to the Eckart-Young-Mirsky theorem, an optimal solution to problem equation 4.3 is given by:
$A^{*}=U_{m}\Sigma_{m}^{1/2},\quad B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top},$ (4.4)
@@
$\|W-A^{*}B^{*}\|_{F}^{2}=\sum_{i=m+1}^{\min(n,mN)}\sigma_{i}^{2}.$ (4.5)
4.2 Minimizing Factorization Residuals
In the previous subsection, we established methods for transforming MoE models into their MoLE counterparts from a theoretical perspective. Here, we address a critical aspect of this transformation: minimizing the residual error that inevitably arises when factorizing expert weights. We begin by establishing the precise conditions under which exact factorization is possible.
Theorem 1.
$K\subseteq\bigcap_{i=1}^{N}\ker(W^{i}).$ (4.6)
Proof.
$W^{i}x=A^{i}Bx=A^{i}\cdot 0=0,$ (4.7)
Sufficiency: Suppose there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ such that $K\subseteq\ker(W^{i})$ for all $i$. We can construct $B\in\mathbb{R}^{m\times n}$ such that $\ker(B)=K$. Since $\dim(K)=n-m$, the matrix $B$ has rank $m$ by the rank-nullity theorem. For each $W^{i}$, the inclusion $K\subseteq\ker(W^{i})$ implies that any vector mapped to zero by $B$ is also mapped to zero by $W^{i}$. By the fundamental theorem of linear algebra, this means $\text{Row}(W^{i})\subseteq\text{Row}(B)$, where $\text{Row}(\cdot)$ denotes the row space. Therefore, there exists $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for each $i$. ∎
Given that exact factorization is generally unattainable, we now consider how to minimize the approximation error through strategic rank reduction. The rank-nullity theorem (Lang, 1987) states that for any linear mapping $W:X\rightarrow Y$:
$\text{rank}(W)+\dim(\ker(W))=\dim(X).$ (4.8)
We implement this approach by computing low-rank approximations of each $W^{i}$ before attempting factorization. Importantly, our empirical experiments in Appendix A demonstrate that this rank reduction has minimal impact on model performance, suggesting that these FFN operators in MoE models inherently possess low-rank structure that can be exploited for more efficient parameterization.
4.3 Transfer MoE to MoLE: A Unified Framework
Based on our theoretical analyses in Sections 4.1 and 4.2, we now propose a unified, systematic framework for transforming Mixture of Experts (MoE) models into their more parameter-efficient Mixture of Latent Experts (MoLE) counterparts. Our framework consists of two principal steps, carefully designed to preserve model capabilities while enabling the latent parameterization.
@@ 4.3 Transfer MoE to MoLE: A Unified Framework
We formalize our approach in Algorithm 1, which provides a complete computational procedure for transforming MoE parameters into the MoLE architecture.
Algorithm 1 Transformation of MoE to MoLE
@@
This algorithm provides a computationally efficient procedure for transforming standard MoE layers into MoLE architecture. The rank reduction parameter $r$ and the latent dimension $m$ serve as hyperparameters that can be tuned to balance performance preservation against parameter efficiency. Typically, we set $r\leq m<\min(n,\sum_{i=1}^{N}m)$, where $n$ is the input dimension.
Our framework is applicable to all linear operators within an MoE layer, including the up, down, and gate operators. By applying this transformation to each operator independently, we convert the entire MoE model to MoLE while maintaining model quality, as demonstrated in our experimental evaluations.
@@ 5 Experiments
This section presents a comprehensive comparative analysis between Mixture of Latent Experts (MoLE) and Mixture of Experts (MoE) architectures in the context of pretraining GPT-2 (Radford et al., 2019) models and evaluating their subsequent performance on downstream tasks.
5.1 Experimental Setup
@@
The MoE and MoLE models were configured with identical architectural parameters except for their Feed-Forward Network (FFN) layer structures, as detailed in Table 6. Both architectures implement $N=32$ experts in their respective FFN layers. For the MoLE model, we establish $k=8$, indicating that each group of $8$ experts shares a common latent representation space. All models were trained on the Wikipedia English (Foundation, ) with sequences truncated to a maximum length of $512$ tokens. Training was conducted using the AdamW optimizer with consistent hyperparameters across all experimental conditions. The comprehensive specifications of model architectures and training configurations are presented in Table 6.
5.2 Parameter Efficiency
@@
As quantified in Table 6, the introduction of shared latent spaces in MoLE architecture yields a substantial reduction in model parameter count compared to the standard MoE architecture. Specifically, MoLE achieves a 40% reduction in non-embedding parameters while maintaining comparable representational capacity. This parameter efficiency represents a significant advancement in model scalability without sacrificing performance.
5.3 Training Dynamics
@@
Figure 2 illustrates the pretraining convergence characteristics of both MoE and MoLE models. The training loss trajectories reveal that MoLE maintains competitive optimization dynamics despite its significantly reduced parameter count. Although the MoE model converges to marginally lower loss values, this difference is negligible when considering the substantial parameter efficiency gained with MoLE. These results suggest that the shared latent representation in MoLE effectively preserves the essential modeling capacity while eliminating redundant parameterization inherent in traditional MoE architectures.
Figure 2: Comparison of training loss curves between MoE and MoLE models on the English Wikipedia dataset. The convergence patterns demonstrate comparable performance, with MoLE exhibiting slightly higher loss values while utilizing substantially fewer parameters than the standard MoE architecture.
@@ 6 Conclusion
In this paper, we presented Mixture of Latent Experts (MoLE), a novel parameter-efficient architecture that addresses the fundamental limitations of traditional MoE models. By factorizing expert weight matrices into shared projections and expert-specific transformations within a lower-dimensional latent space, MoLE significantly reduces parameter count and computational overhead without sacrificing model performance. Our comprehensive empirical evaluation demonstrates that MoLE not only preserves the capabilities of standard MoE models across diverse language tasks but also enables more economical scaling of large language models.
The theoretical framework we developed for transforming pre-trained MoE models into the MoLE architecture provides a principled approach to expert parameterization, offering insights into the nature of redundancy in neural networks. As models continue to grow in scale and complexity, architectures like MoLE that prioritize parameter efficiency while maintaining performance will become increasingly valuable for practical deployment scenarios. Future work could explore extending these factorization techniques to other components of the transformer architecture and investigating dynamic adaptation of the latent space dimensionality based on input complexity.
Appendix A Parameter Redundancy in Neural Networks
This section provides empirical evidence for significant parameter redundancy within Feed-Forward Network (FFN) layers, thereby substantiating the theoretical framework presented in Section 4. We conduct our analysis using Qwen1.5-MoE-A2.7B [Team, 2024], a state-of-the-art Mixture-of-Experts architecture comprising $14.3$ billion total parameters while activating only $2.7$ billion parameters during inference. Each FFN layer in this architecture is characterized by an intermediate dimension $m=1408$, hidden dimension $n=2048$, and $N=60$ distinct expert modules.
Figure 3: Singular value distributions of expert modules in the Qwen1.5-MoE-A2.7B model. The relatively uniform distributions indicate that weight matrices exhibit high-rank characteristics, with no significant concentration of information in a small subset of singular values.
As illustrated in Figure 3, the singular value distributions of the expert modules demonstrate an absence of dominant singular values—a finding that contradicts conventional assumptions about the inherent low-rank structure of neural network weight matrices. This observation suggests that naive dimensional reduction techniques applying uniform low-rank approximations would potentially eliminate performance-critical parameters.
Nevertheless, we demonstrate that despite the apparent full-rank nature of the weight matrices $W_{\text{up}}$, $W_{\text{gate}}$, and $W_{\text{down}}$, this characteristic is not a prerequisite for optimal model performance. Through systematic experimentation, we establish that these operators can be effectively replaced with carefully calibrated low-rank approximations while maintaining comparable performance metrics.
For our analysis, we define a ratio-$r$ low-rank approximation $\tilde{W}$ of any matrix $W$ as a matrix whose rank equals $r\times\text{rank}(W)$, where $0<r\leq 1$. In accordance with the Eckart-Young-Mirsky theorem [Schmidt, 1989], these approximations are computed via Singular Value Decomposition (SVD), retaining only the largest $r\times\text{rank}(W)$ singular values and their corresponding singular vectors.
To rigorously assess model capabilities under low-rank constraints, we evaluate performance on three benchmark tasks: MMLU [Hendrycks et al., 2021], GSM8K [Cobbe et al., 2021], and Wikitext-2 [Merity et al., 2016]—the same evaluation suite utilized in the original Qwen report. All experiments were conducted using the lm-evaluation-harness evaluation framework [Gao et al., 2024]. Table 2 presents the comparative results.
$1.0$ (Original) 60.1 61.0 9.49
$0.8$ 61.2 60.3 9.65
$0.6$ 60.1 59.4 9.85
Table 2: Performance comparison of QwenMoE models with varying low-rank approximation ratios across multiple benchmarks. For the low-rank ratio $r=0.8$ case, the low-rank estimation model performs competitive with the original model on various tasks.
Table 2 illustrates the relationship between rank reduction and model performance. The baseline case ($r=1.0$) represents the original, unmodified model with full-rank weight matrices. Notably, when reducing the rank of FFN operators by 20% ($r=0.8$), we observe no significant performance degradation. In fact, the reduced-rank model demonstrates superior performance on the GSM8K benchmark, achieving a 1.1 percentage point improvement over the original model, while maintaining comparable performance on MMLU accuracy and Wikitext-2 PPL.
These empirical findings provide compelling evidence that, despite the mathematical full-rank property of FFN weight matrices, a substantial proportion of parameters contain redundant information that can be effectively approximated through lower-dimensional representations. This parameter redundancy phenomenon forms the empirical foundation for our theoretical analysis in Section 4.
Appendix B Transformation from MoE to MoLE
This section presents our empirical analysis of transforming standard MoE architectures into their corresponding MoLE counterparts. We specifically investigate the Qwen1.5-MoE-A2.7B model and systematically examine the impact of varying the latent space configuration parameter $k$ on model performance.
As established in Section 3, the parameter $k$ in MoLE architecture represents a critical hyperparameter that controls the trade-off between parameter efficiency and model expressivity. We evaluate three distinct configurations of $k\in\{1,10,60\}$, where:
  • $k=1$ corresponds to the original MoE architecture (baseline)
  • $k=10$ represents a balanced MoLE configuration with multiple latent spaces
  • $k=60$ represents the extreme case with a single shared latent space
For consistent evaluation with Section A, we assess model performance on three established benchmarks: MMLU (reasoning), GSM8K (mathematical reasoning), and Wikitext-2 (language modeling perplexity).
Table 3 presents the comparative performance across different MoLE configurations.
Number of experts per latent ($k$) GSM8K (%) $\uparrow$ MMLU (%) $\uparrow$ Wikitext-2 PPL $\downarrow$
$1$ (Original MoE) 60.1 61.0 9.49
$10$ 59.4 60.0 9.74
$60$ 14.3 40.1 25.97
Table 3: Performance comparison of Qwen1.5-MoE-A2.7B with varying MoLE configurations. The balanced configuration ($k=10$) maintains near-parity with the original model while achieving significant parameter reduction. The single latent space configuration ($k=60$) demonstrates substantial performance degradation.
The results demonstrate a clear relationship between latent space configuration and model performance. The balanced MoLE configuration ($k=10$) achieves comparable performance to the original MoE architecture, with only marginal degradation across all benchmarks: a 0.7 percentage point decrease on GSM8K, 1.0 percentage point decrease on MMLU, and 2.6% increase in perplexity on Wikitext-2. This configuration successfully reduces parameter count while preserving most of the model’s capabilities.
In contrast, the extreme configuration ($k=60$) with a single shared latent space per FFN layer exhibits catastrophic performance degradation: a 45.8 percentage point decrease on GSM8K, 20.9 percentage point decrease on MMLU, and 173.7% increase in perplexity. This dramatic performance collapse empirically validates our theoretical analysis that multiple latent spaces are necessary to preserve the expressive capacity of the original MoE architecture.
These findings provide compelling evidence that the MoLE architecture, when properly configured with an appropriate number of latent spaces, can achieve substantial parameter efficiency while maintaining competitive performance. Furthermore, they establish clear boundaries for the practical application of latent space factorization in expert-based neural architectures.
Appendix C Critical Role of the “Down Operator”

### v2

@@ 5 Transformation from MoE to MoLAE
In this section, we establish the theoretical foundation for transforming a standard MoE model into its corresponding MoLAE counterpart. We focus on the “up operator” as a representative example. For analytical clarity, we make two simplifying assumptions: (1) we omit the subscript of $W_{\text{up}}^{i}$ for ease of notation during this analysis, and (2) we consider the case where $k=N$, implying a single shared latent space operator for all experts.
Keep the “up operator”. Experiments on Qwen1.5-MoE-A2.7B indicate that transferring all operators during MoLAE transformation does not hurt performance. However, for Moonlight-16B-A3B, empirical results (Appendix B) show superior performance when preserving the “up operator” structure. Thus, for Moonlight, only its “down” and “gate” operators are converted to the MoLAE structure, retaining the “up operator”. Conversely, all operators are transferred for Qwen models.
@@ 5.1 Transformation via Matrix Factorization
For a given weight matrix $W^{i}$ associated with expert $i$, we aim to find corresponding matrices $A^{i}$ and $B$ such that $W^{i}X\approx A^{i}BX$ for the activation $X$. This transformation represents the post-training perspective where we seek to convert pre-trained MoE parameters into the MoLAE architecture.
@@
$\min_{A^{i},B}\quad F(A^{1},\cdots,A^{N},B):=\frac{1}{2}\sum_{i=1}^{N}\|W^{i}-A^{i}B\|_{F}^{2},$ (6)
where $\|\cdot\|_{F}$ denotes the Frobenius norm. Problem (6) admits the theoretical optimal solutions by using the SVD decomposition technique.
Closed-Form Solution via Singular Value Decomposition. Obviously, problem (6) has infinitely many solutions since one can change one optimal solution $(A^{*},B^{*})$ into another by selecting a constant $\lambda$ and obtain another optimal solution $(\lambda A^{*},\frac{1}{\lambda}B^{*})$. Hence, we only provide “one” optimal solution here.
To derive a closed-form solution to problem (6), we consolidate the matrices into the following block structures:
$W=\begin{pmatrix}W^{1}\\ W^{2}\\ \vdots\\ W^{N}\end{pmatrix},\quad A=\begin{pmatrix}A^{1}\\ A^{2}\\ \vdots\\ A^{N}\end{pmatrix}.$ (7)
With this notation, problem (6) can be reformulated as:
$\min_{A,B}\quad\frac{1}{2}\|W-AB\|_{F}^{2},$ (8)
@@
While problem (8) has infinitely many solutions due to its underdetermined nature, the Eckart-Young-Mirsky theorem (Schmidt, 1989) provides an optimal solution with respect to the Frobenius norm. Specifically, let $W=U\Sigma V^{\top}$ be the singular value decomposition (SVD) of $W$, where:
@@
Let $\Sigma_{m}$ be the truncated version of $\Sigma$ that retains only the $m$ largest singular values: $\Sigma_{m}=\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m},0,\ldots,0)\in\mathbb{R}^{mN\times n}$. According to the Eckart-Young-Mirsky theorem, an optimal solution to problem (8) is given by:
$A^{*}=U_{m}\Sigma_{m}^{1/2},\quad B^{*}=\Sigma_{m}^{1/2}V_{m}^{\top},$ (9)
@@
$\|W-A^{*}B^{*}\|_{F}^{2}=\sum_{i=m+1}^{\min(n,mN)}\sigma_{i}^{2}.$ (10)
To ensure the effectiveness of this transformation, we provide more details in Appendix A which discuss how to minimize the factorization residuals during the transformation.
5.2 Transfer MoE to MoLAE: A Unified Framework
Based on our theoretical analyses above, we propose a unified, systematic framework for transforming Mixture of Experts (MoE) models into their more parameter-efficient Mixture of Latent Experts (MoLAE) counterparts. Our framework consists of two principal steps, carefully designed to preserve model capabilities while enabling the latent parameterization.
@@ 5.2 Transfer MoE to MoLAE: A Unified Framework
We formalize our approach in Algorithm 1, which provides a complete computational procedure for transforming MoE parameters into the MoLAE architecture.
Algorithm 1 Transformation of MoE to MoLAE
@@
This algorithm provides a computationally efficient procedure for transforming standard MoE layers into MoLAE architecture. The rank reduction parameter $r$ and the latent dimension $m$ serve as hyperparameters that can be tuned to balance performance preservation against parameter efficiency. Typically, we set $r\leq m<\min(n,\sum_{i=1}^{N}m)$, where $n$ is the input dimension.
Our framework is applicable to all linear operators within an MoE layer, including the up, down, and gate operators. By applying this transformation to each operator independently, we can convert the MoE model to MoLAE while maintaining overall performance.
@@ 6 Experiments
In this section, we evaluate the effectiveness of MoLAE on downstream tasks and the pre-training performance on GPT-2 (Radford et al., 2019). More experiments, like the importance of different “up, gate, down” operators and model configurations, are provided in Appendix B.
Table 3: Comparisons of Qwen1.5-MoE-A2.7B and Moonlight-16B-A3B to their MoLAE architectures. To simplify the computations of efficiency, we only consider the MoE parts of the model.
Models Benchmarks Params
MMLU $\uparrow$ GSM8K $\uparrow$ CEval $\uparrow$ MNLI $\uparrow$ Wiki PPL $\downarrow$
Qwen1.5-MoE 61.0 60.1 80.2 49.9 9.49 14.3B
Qwen1.5-MoLAE
($k$=10)
60.0 59.4 74.1 40.6 9.74 10.5B
[…]
Moonlight-MoLAE
($k$=8)
60.2 72.3 62.8 38.7 10.6 12.3B
6.1 Transformation from MoE to MoLAE: Downstream Tasks
We first present our empirical analysis of transforming standard MoE architectures into their corresponding MoLAE counterparts. We specifically investigate two popular MoE models, including Qwen1.5-MoE-A2.7B model Team (2024) and Moonlight-16B-A3B (Liu et al., 2025), on diverse tasks, such as MMLU (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), CEval (Huang et al., 2023), MNLI (Wang et al., 2019) and Wikitext-2 (Merity et al., 2016).
As shown in Table 3, MoLAE achieves performance comparable to that of the original MoE models, while exhibiting notable parameter efficiency and only slight performance degradations across the benchmarks. This result highlights the effectiveness of the transformation from MoE to MoLAE architectures. Furthermore, as indicated in Section 4, the group size $k$ in MoLAE architecture represents a critical hyperparameter that controls the trade-off between parameter efficiency and model expressivity. We evaluate several distinct configurations of $k\in\{1,10,20,30,60\}$ on the Qwen1.5-MoE model, which consists of 60 experts. Specifically, $k=1$ corresponds to the original MoE architecture; $k=10$ and $k=20$ represent balanced MoLAE configurations with multiple latent spaces; while $k=30$ and $k=60$ denote extreme cases with only one or two shared latent spaces, respectively. As shown in Table 2, with a moderate group size, such as $k=10$ for the Qwen1.5-MoE model, MoLAE largely preserves performance across various benchmarks. However, as the group size increases, the performance of the MoLAE model progressively deteriorates. For simpler tasks, such as CEval and MNLI, the performance decline is minimal, indicating that MoLAE retains most of the model’s capacity. In contrast, for more challenging tasks, such as GSM8K, there is a significant performance degradation. This dramatic decline empirically supports our theoretical analysis, which suggests that multiple latent spaces are essential to maintain the capacity of the original MoE model.
Figure 2: Ablation study of group size $k$ on the Qwen1.5-MoE model.
Figure 3: Comparison of training loss curves between MoE and MoLAE models on the English Wikipedia dataset.
6.2 Pretraining of MoLAE
@@
To further validate the effectiveness of MoLAE, we construct the paired MoE and MoLAE models derived from the GPT-2 model (Radford et al., 2019), and then conduct the pretraining on the models. The MoE (151M) and MoLAE (94M) models are configured with identical architectural parameters except for FFN layer structures, as detailed in Table 5. Both architectures implement $N=32$ experts in their respective FFN layers. For MoLAE model, we establish $k=8$, indicating that each group of $8$ experts shares a common latent representation space. All models are trained on the Wikipedia English (Foundation, ) with maximum length of $512$. Models are updated using AdamW optimizer with consistent hyper-parameters across all runs.
Parameter Efficiency
@@
As quantified in Table 5, the introduction of shared latent spaces in MoLAE architecture yields a substantial reduction in model parameter count compared to the standard MoE architecture. Specifically, MoLAE achieves a 40% reduction in non-embedding parameters while maintaining comparable model capacity. This parameter efficiency represents a significant advancement in model scalability without sacrificing performance.
Training Dynamics
@@
Figure 3 illustrates the pretraining convergence characteristics of both MoE and MoLAE models. The training loss trajectories reveal that MoLAE maintains competitive optimization dynamics despite its significantly reduced parameter count. Although the MoE model converges to marginally lower loss values, this difference is negligible when considering the substantial parameter efficiency gained with MoLAE. These results suggest that the shared latent representation in MoLAE effectively preserves the essential modeling capacity while eliminating redundant parameterization inherent in traditional MoE architectures. Using the well-trained MoE and MoLAE models, we evaluated their performance on the downstream task of Wikitext-2 perplexity. The results indicate that the performance gap between the two models is minimal, with PPL values of 79.5 and 81.5, respectively. This suggests that the MoLAE architecture serves as an effective pretraining base model, offering excellent efficiency while maintaining competitive performance.
@@ 7 Conclusion
We introduce Mixture of Latent Experts (MoLAE), which overcomes limitations of traditional MoE models by factorizing expert weights into shared projections and expert-specific transformations in a lower-dimensional space. This approach reduces parameters and computation while maintaining performance across language tasks. Our theoretical framework for converting pre-trained MoE models to MoLAE provides insights into neural network redundancy. As models grow, such parameter-efficient architectures become increasingly valuable. Future work could extend these techniques to other transformer components and explore dynamic latent space adaptation.
Appendix A Minimizing Factorization Residuals
In the previous subsection 5, we established methods for transforming MoE models into their MoLAE counterparts from a theoretical perspective. Here, we address a critical aspect of this transformation: minimizing the residual error that inevitably arises when factorizing expert weights. We begin by establishing the precise conditions under which exact factorization is possible.
Theorem 1
$K\subseteq\bigcap_{i=1}^{N}\ker(W^{i}).$ (11)
Proof 1
$W^{i}x=A^{i}Bx=A^{i}\cdot 0=0,$ (12)
Sufficiency: Suppose there exists an $(n-m)$-dimensional subspace $K\subseteq\mathbb{R}^{n}$ such that $K\subseteq\ker(W^{i})$ for all $i$. We can construct $B\in\mathbb{R}^{m\times n}$ such that $\ker(B)=K$. Since $\dim(K)=n-m$, the matrix $B$ has rank $m$ by the rank-nullity theorem. For each $W^{i}$, the inclusion $K\subseteq\ker(W^{i})$ implies that any vector mapped to zero by $B$ is also mapped to zero by $W^{i}$. By the fundamental theorem of linear algebra, this means $\text{Row}(W^{i})\subseteq\text{Row}(B)$, where $\text{Row}(\cdot)$ denotes the row space. Therefore, there exists $A^{i}\in\mathbb{R}^{m\times m}$ such that $W^{i}=A^{i}B$ for each $i$.
Given that exact factorization is generally unattainable, we now consider how to minimize the approximation error through strategic rank reduction. The rank-nullity theorem [Lang, 1987] states that for any linear mapping $W:X\rightarrow Y$:
$\text{rank}(W)+\dim(\ker(W))=\dim(X).$ (13)
We implement this approach by computing low-rank approximations of each $W^{i}$ before attempting factorization. Importantly, our empirical experiments in Appendix 3.2 demonstrate that this rank reduction has minimal impact on model performance, suggesting that these FFN operators in MoE models inherently possess low-rank structure that can be exploited for more efficient parameterization.
Appendix B Critical Role of the “Up Operator”